If segment XY has a Midpoint M, use X (2 , 10) and M (-3 , 15) to find the missing endpoint Y.
Answer by stanbon(75887)

If segment XY has a Midpoint M, use X (2 , 10) and M (-3 , 15) to find the missing endpoint Y.
-------
Let the coordinates of "Y" be (x,y)
-----
Since M is the midpoint,
(x+2)/2 = -3
x+2 = -6
x = -8
-----------------
(y+10)/2 = 15
y+10 = 30
y = 20
----------------
Cheers,
Stan H.
-----------